OpenStudy (anonymous):

the area of a rectangle is 77 ft. squared and the length is 3 ft more than twice the width. Find the dimensions

OpenStudy (neer2890):

Area of rectangle= length* breadth 77=(3+2w)*w find w= breadth length=2+3w

OpenStudy (neer2890):

\[2w ^{2}+3w-77=0\] it's quadratic equation, use this formula for finding roots of quadratic equation\[ax ^{2}+bx+c=0\] \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

I think it will factor, if you are interested in less computation.